OpenStudy (anonymous):

Find the surface area of a sphere with a diameter of 14 cm. Leave your answer in terms of pi @Directrix

Directrix (directrix):

sphere, volume, radius = 7 cm, Radius is half of diameter. 14/2 = 7 Volume of Sphere: V = 4/3 * π * r³

Directrix (directrix):

V = 4/3 * π * r³ V = 4/3 * π * 7³

Directrix (directrix):

In terms of π, remember.

OpenStudy (anonymous):

i think i did it wrong

OpenStudy (anonymous):

i got 343 and i didnt approximate the pi

Directrix (directrix):

I got (1372/3) π cubic cm

Directrix (directrix):

I think you cubed the 7 and did not multiply that times (4/3). Try again.

OpenStudy (anonymous):

yea i cubed it

Directrix (directrix):

What are the answer options? I don't think we are supposed to approximate 1372/3

OpenStudy (anonymous):

457.3?

Directrix (directrix):

Answer options?

Directrix (directrix):

There has to be a pi.

OpenStudy (anonymous):

A.1,219 m2 B.50,265 m2 C.181 m2 D1,307 m2

Directrix (directrix):

They approximated the improper fraction. Let me think.

Directrix (directrix):

(1372/3) π cubic cm = 457.3 π cubic meters. HOLD ON: We did volume rather than Surface Area. So, we have the right answer to the wrong question.

Directrix (directrix):

The answer we got is an option because the homework writer always includes the volume and surface area answers hoping that we will mess up and we did.

Directrix (directrix):

No worries. We will fix it.

OpenStudy (anonymous):

sorry i put the wrong answer choices

OpenStudy (anonymous):

196pi cm^2

Directrix (directrix):

I put the wrong formula.

Directrix (directrix):

Surface area of a sphere: A = 4 * π * r²

Directrix (directrix):

A = 4 * π * 7²

Directrix (directrix):

I got 196π square cm.

Directrix (directrix):

What did you get?

OpenStudy (anonymous):

the same

Directrix (directrix):

Is that one of the answer options?

OpenStudy (anonymous):

yes

Directrix (directrix):

Sorry that I messed us up on that question.

OpenStudy (anonymous):

no that was my fault but you did it before so that's how i got the answer